# GraphRCG: Self-conditioned Graph Generation via Bootstrapped Representations

### v1

@@ Abstract.
Graph generation generally aims to create new graphs that closely align with a specific graph distribution. Existing works often implicitly capture this distribution through the optimization of generators, potentially overlooking the intricacies of the distribution itself. Furthermore, these approaches generally neglect the insights offered by the learned distribution for graph generation. In contrast, in this work, we propose a novel self-conditioned graph generation framework designed to explicitly model graph distributions and employ these distributions to guide the generation process. We first perform self-conditioned modeling to capture the graph distributions by transforming each graph sample into a low-dimensional representation and optimizing a representation generator to create new representations reflective of the learned distribution. Subsequently, we leverage these bootstrapped representations as self-conditioned guidance for the generation process, thereby facilitating the generation of graphs that more accurately reflect the learned distributions. We conduct extensive experiments on generic and molecular graph datasets across various fields. Our framework demonstrates superior performance over existing state-of-the-art graph generation methods in terms of graph quality and fidelity to training data.
copyright: acmcopyrightconference: SIGKDD; August 25-29, 2024; Barcelona, Spain
1. Introduction
@@
The task of generating graphs that align with a specific distribution plays a crucial role in various fields such as drug discovery (Shi et al., 2019), public health (Guo et al., 2021), and traffic modeling (Yu and Gu, 2019). In recent times, deep generative models have been prevalently studied to address the problem of graph generation. Unlike conventional methods that rely on random graph models, recent methods generally learn graph distributions through advanced deep generative models, e.g., variational autoencoders (VAEs) (Guo et al., 2020; Wang et al., 2022), generative adversarial networks (GANs) (Gamage et al., 2020; De Cao and Kipf, 2018), normalizing flows (Zang and Wang, 2020; Luo et al., 2021), and diffusion models (Lee et al., 2023; Niu et al., 2020; Vignac et al., 2022). These models excel at capturing complex structural patterns in graphs, enabling the creation of new graphs with desirable characteristics.
Figure 1. The comparison between the conventional framework and ours. Instead of directly learning the graph distribution, we encode graphs into representations, and learn their distributions for further utilization during generation.
@@
Despite these advances, the precise modeling and utilization of graph distributions, although crucial for high-fidelity generation, remains underexplored. In fact, it is essential to accurately capture and utilize important patterns in the training data for generation (Kong et al., 2023; Karami, 2023), particularly in complex scenarios like molecular graph generation (Du et al., 2022). For example, precise modeling of molecular properties is key to optimizing molecular structures while maintaining similarity to known molecules. However, the prevalent strategy is to use reconstruction loss to implicitly embed graph distribution within the generator, which may compromise effectiveness. Moreover, the utilization of the captured graph distributions is also less investigated. Ideally, generators should be designed to explicitly guide the generative process, ensuring that the output graphs closely follow the defined graph distributions. Nonetheless, existing research tends to rely on simple features to control generation, such as molecular characteristics (Vignac et al., 2022) or degree information (Chen et al., 2023). Such a strategy requires domain knowledge to design the specific features, while also lacking more comprehensive modeling of the entire distribution. Therefore, the study of graph generation is confronted with two crucial research questions (RQs): (RQ1) Capturing Distributions. How to precisely capture the graph distribution with rich information helpful for the generation process? (RQ2) Utilizing Distributions. How to adeptly harness these distributions as direct guidance for the generation of graphs? Addressing these challenges is essential for high-fidelity graph generation.
@@
In practice, however, the above research questions present significant challenges due to the intricate nature of graph data. (1) Complex Dataset Patterns. Real-world graphs, such as social networks (Chen et al., 2023) and molecular structures (Lee et al., 2023), exhibit highly complex patterns. These include varying degrees of sparsity, inconsistent clustering coefficients, and specific distributions of node and edge attributes (Huang et al., 2022). Capturing these complex patterns accurately through generative models can be particularly challenging. (2) Discrete Sequential Generation. Unlike images, where generation is often treated as a pixel-wise or patch-wise process (Ho et al., 2022; Rombach et al., 2022; Dhariwal and Nichol, 2021), the generation of graphs is inherently sequential and discrete (You et al., 2018b; Wang et al., 2022; Niu et al., 2020). That being said, graphs are generated through a sequence of steps, each with significant implications, such as modifying chemical properties via the addition or removal of atoms and bonds (Shi et al., 2019; Kong et al., 2023; Karami, 2023). As a result, it is suboptimal to directly guide generation toward true distributions, particularly in the initial steps of the process with graphs largely deviating from the learned distribution.
@@
  • (1) In this work, we explore the potential and importance of capturing and utilizing training data distributions for graph generation to enhance generation performance.
@@
  • (2) We innovatively propose a self-conditioned graph generation framework to capture and utilize training data distributions via bootstrapped representations with our devised self-conditioned modeling and self-conditioned guidance, respectively.
@@
  • (3) We perform a systematic study to evaluate the performance of our framework in a variety of real-world and synthetic datasets. The results demonstrate the effectiveness of our framework in comparison to other state-of-the-art baselines.
2. Related Works
2.1. Denoising Diffusion Models for Generation
@@
Recently, denoising diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) have been widely adopted in various fields of generation tasks. Their notable capabilities span various domains, including image generation (Dhariwal and Nichol, 2021; Rombach et al., 2022; Ho et al., 2022), text generation (Austin et al., 2021; Li et al., 2022; Chen et al., 2022), and temporal data modeling (Tashiro et al., 2021; Lopez Alcaraz and Strodthoff, 2023). In general, denoising diffusion models are probabilistic models designed to learn a data distribution by denoising variables with normal distributions (Song et al., 2020; Kong and Ping, 2021). In particular, they first create noisy data by progressively adding and intensifying the noise in the clean data, in a Markovian manner. Subsequently, these models learn a denoising network to backtrack each step of the perturbation. During training, the denoising networks are required to predict the clean data or the added noise, given the noisy data after perturbation. After optimization, the denoising networks could be used to generate new data via iterative denoising of noise sampled from a prior distribution (San-Roman et al., 2021; Vahdat et al., 2021).
2.2. Graph Generation
@@
Based on the strategies used, graph generation methods could be classified into two categories: (1) One-shot Generation. In this category, the models are required to generate all edges among a defined node set in one single step. One-shot generation models are typically built upon the Variational Autoencoder (VAE) or the Generative Adversarial Network (GAN) structure, aiming to generate edges independently based on the learned latent embeddings. Normalizing flow models (Zang and Wang, 2020; Luo et al., 2021) propose to estimate the graph density, by establishing an invertible and deterministic function to map latent embeddings to the graphs. More recently, diffusion models have also been adopted for graph generation (Lee et al., 2023; Jo et al., 2022). To deal with the discrete nature of graph data, DiGress (Vignac et al., 2022) proposes to leverage discrete diffusion by considering node and edge types as states in the Markovian transition matrix. (2) Sequential Generation. This strategy entails generating graphs through a series of sequential steps, typically by incrementally adding nodes and connecting them with edges. Models in this category often utilize recurrent networks (Li et al., 2018; You et al., 2018b) or Reinforcement Learning (RL) (You et al., 2018a) to guide the generation process (Shi et al., 2019; Ahn et al., 2021). Sequential generation is particularly suitable for generating graphs with specific desired properties (Zhu et al., 2022).
2.3. Conditional Generation
@@
Recent works have also explored various strategies to condition generation on specific classes or features. For example, ARM (Dhariwal and Nichol, 2021) proposes to utilize gradients from classifiers to guide the generation process within each step. LDM (Rombach et al., 2022) enables the incorporation of external information, such as text (Reed et al., 2016) and semantic maps (Isola et al., 2017), with a specific encoder to learn the representations. The representations are then incorporated into the underlying UNet backbone (Ronneberger et al., 2015). In RCG (Li et al., 2023), the authors employ a self-conditioned strategy to condition generation on representations learned from a pre-trained encoder. Despite the advancements in image generation, it still presents significant difficulty when applying these methods in graph generation, due to the complex dataset patterns (Chen et al., 2023; Lee et al., 2023; Huang et al., 2022) and discrete sequential generation (Wang et al., 2022; Niu et al., 2020; Shi et al., 2019). In contrast, our framework deals with these challenges with self-conditioned modeling and guidance to enhance graph generation efficacy.
@@
Figure 2. The overall process of our framework during training. Specifically, we learn from the training data a representation generator that outputs a representation based on noise sampled from a standard Gaussian distribution. After that, a graph generator is trained to generate new graphs from noisy graphs under the guidance of the representations.
3. Self-Conditioned Graph Generation
@@
Our self-conditioned graph generation framework consists of two modules: self-conditioned modeling and self-conditioned guidance for RQ1 and RQ2, respectively. As illustrated in Fig. 2, in self-conditioned modeling, we employ a representation generator to capture graph distributions by learning to denoise representations with noise. The alignment loss between noisy representations and noisy graphs acts as a self-supervised loss to train the encoder. With the optimized representation generator, we train the graph generator by denoising graphs with added noise. At each generation step, we perform self-conditioned guidance via bootstrapped representations with noise at the same timestep. We further enhance guidance performance by aligning the denoised graph with the clean representations.
@@
3.1. Self-conditioned Modeling
@@
Representation Generator. Given the representations of all training data provided by the graph encoder, the representation generator is required to learn to generate representations with the same distribution. To enhance generation performance, following (Li et al., 2023), we utilize the Representation Diffusion Model (RDM) architecture, which generates representations from Gaussian noise, based on the process of Denoising Diffusion Implicit Models (DDIM) (Song et al., 2021). In particular, RDM utilizes a backbone comprising a fully connected network with multiple residual blocks. Each block is composed of an input layer, a timestep embedding projection layer, and an output layer. The number of residual blocks and the hidden dimension size both act as hyper-parameters. During training, given a representation of a graph sample, learned by the graph encoder, we first perturb it by adding random noise as follows:
(1) $\mathbf{h}_{t}=\sqrt{\alpha_{t}}\mathbf{h}_{0}+\sqrt{1-\alpha_{t}}\epsilon,\quad\text{where}\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$
@@
(2) $\mathcal{L}_{RG}=\mathbb{E}_{\mathbf{h}_{0},\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\mathbf{h}_{0}-f_{\gamma}\left(\mathbf{h}_{t},t\right)\right\|_{2}^{2}\right],$
where $\mathbf{h}_{0}$ is sampled from representations of graphs in training data, and $t$ is uniformly sampled from $\{1,2,\dotsc,T\}$. The target of the representation generator $f_{\gamma}$ is to capture the representation distribution and learn to generate representations from random noise. After optimization, the representation generator could perform sampling from random noise, following the DDIM strategy (Song et al., 2021), to obtain new representations.
@@
(3) $\mathcal{L}_{AR}=\mathbb{E}_{G,t}\left[\left\|\mathbf{h}_{t}-h_{\eta}(\mathbf{X}_{t},\mathbf{E}_{t})\right\|_{2}^{2}\right],$
@@
(4) $q(G_{t}|G_{t-1})=(\mathbf{X}_{t-1}\mathbf{Q}_{t}^{\mathbf{X}},\mathbf{E}_{t-1}\mathbf{Q}_{t}^{\mathbf{E}}),$
@@
(5) $q(G_{t}|G_{0})=(\mathbf{X}_{0}\prod_{i=1}^{t}\mathbf{Q}_{i}^{\mathbf{X}},\mathbf{E}_{0}\prod_{i=1}^{t}\mathbf{Q}_{i}^{\mathbf{E}}),$
@@
To specify the Markov transition matrices, previous works have explored several feasible choices. The most prevalent choices in the literature have been uniform transitions (Austin et al., 2021; Yang et al., 2023) and absorbing transitions (Chen et al., 2023; Kong et al., 2023). However, these do not contain the distribution information and thus could not benefit the capturing of graph distribution in our framework. Therefore, in our approach, we leverage the marginal transitions (Ingraham et al., 2023; Vignac et al., 2022), in which the probability of transitioning to any given state is directly related to its marginal probability observed in the dataset. In this manner, the transition matrices are modeled in a way that mirrors the natural distribution of states in graph data. As the edges are generally sparse in the graph data, the probability of jumping to the state of “no edge” is significantly higher than that of other states. To present the noise process, we first define $\mathbf{p}^{\mathbf{X}}\in\mathbb{R}^{a}$ and $\mathbf{p}^{\mathbf{E}}\in\mathbb{R}^{b}$ as the marginal distributions for the node and edge types, respectively. The marginal transition matrices for nodes and edges are formulated as
(6) $\mathbf{Q}_{t}^{\mathbf{X}}=\alpha^{t}\mathbf{I}+\beta^{t}\mathbf{1}_{a}(\mathbf{p}^{\mathbf{X}})^{\top},\ \ \text{and}\ \ \mathbf{Q}_{t}^{\mathbf{E}}$ $=\alpha^{t}\mathbf{I}+\beta^{t}\mathbf{1}_{b}(\mathbf{p}^{\mathbf{E}})^{\top}.$
@@
3.2. Self-conditioned Guidance (Training)
In self-conditioned guidance for RQ2, we optimize a graph generator to create new graphs conditioned on bootstrapped representations from the representation generator. Existing works have explored various methods to perform generation conditioned on specific properties, such as high activity in molecular graph generation (Huang et al., 2022). However, these properties are generally several scalar values, which could capture only a small fraction of the information in the dataset. Moreover, these approaches often adopt a classifier or regressor to guide the generation process, which could not exploit the useful information in representations. In contrast, using representations as guidance for generation could largely benefit from the learned distribution in our representation generator.

### v2

@@ Abstract
Graph generation aims to create new graphs that closely align with a target graph distribution. Existing works often implicitly capture this distribution by aligning the output of a generator with each training sample. As such, the overview of the entire distribution is not explicitly captured and used for graph generation. In contrast, in this work, we propose a novel self-conditioned graph generation framework designed to explicitly model graph distributions and employ these distributions to guide the generation process. We first perform self-conditioned modeling to capture the graph distributions by transforming each graph sample into a low-dimensional representation and optimizing a representation generator to create new representations reflective of the learned distribution. Subsequently, we leverage these bootstrapped representations as self-conditioned guidance for the generation process, thereby facilitating the generation of graphs that more accurately reflect the learned distributions. We conduct extensive experiments on generic and molecular graph datasets. Our framework demonstrates superior performance over existing state-of-the-art graph generation methods in terms of graph quality and fidelity to training data.
1 Introduction
@@
The task of generating graphs that align with a specific distribution plays a crucial role in various fields such as drug discovery [43], public health [13], and traffic modeling [58]. In recent times, deep generative models have been prevalently studied to address the problem of graph generation. Unlike conventional methods that rely on random graph models, recent methods generally learn graph distributions through advanced deep generative models, e.g., variational autoencoders (VAEs) [14, 53], generative adversarial networks (GANs) [10, 6], normalizing flows [59, 32], and diffusion models [26, 34, 52]. These models excel at capturing complex structural patterns in graphs, enabling the creation of new graphs with desirable characteristics.
@@
Despite these advances, the precise modeling and utilization of graph distributions, although crucial for high-fidelity generation, remains underexplored. In fact, it is essential to accurately capture and utilize important patterns in the training data for generation [24, 22], particularly in complex scenarios like molecular graph generation [8]. For example, precise modeling of molecular properties is key to optimizing molecular structures while maintaining similarity to known molecules. However, the prevalent strategy is to use reconstruction loss to implicitly embed graph distribution within the generator, which may compromise effectiveness. Moreover, the utilization of the captured graph distributions is also less investigated. Ideally, generators should be designed to explicitly guide the generative process, ensuring that the output graphs closely follow the defined graph distributions. Nonetheless, existing research tends to rely on simple features to control generation, such as molecular characteristics [52] or degree information [4]. Such a strategy requires domain knowledge to design the specific features, while also lacking more comprehensive modeling of the entire distribution. Therefore, the study of graph generation is confronted with two crucial research questions (RQs): (RQ1) Capturing Distributions. How to precisely capture the graph distribution with rich information helpful for the generation process? (RQ2) Utilizing Distributions. How to adeptly harness these distributions as direct guidance for the generation of graphs? Addressing these challenges is essential for high-fidelity graph generation.
Figure 1: The comparison between the conventional framework and ours. Instead of directly learning the graph distribution, we encode graphs into representations, and learn their distributions for further utilization during generation.
@@
In practice, however, the above research questions present significant challenges due to the intricate nature of graph data. (1) Complex Dataset Patterns. Real-world graphs, such as social networks [4] and molecular structures [26], exhibit highly complex patterns. These include varying degrees of sparsity, inconsistent clustering coefficients, and specific distributions of node and edge attributes [17]. Capturing these complex patterns accurately through generative models can be particularly challenging. (2) Progressive Alignment to Training data. Unlike images, where generation is often treated as a pixel-wise or patch-wise process [16, 38, 7], the generation of graphs is inherently sequential [57, 53, 34]. That being said, graphs are generated through a sequence of steps, each with significant implications, such as modifying chemical properties via the addition or removal of atoms and bonds [43, 24, 22]. As a result, it is suboptimal to directly guide generation toward true distributions, particularly in the initial steps of the process with graphs largely deviating from the learned distribution.
@@
  • In this work, we explore the potential and importance of capturing and utilizing training data distributions for graph generation to enhance generation performance.
@@
  • We innovatively propose a self-conditioned graph generation framework to capture and utilize training data distributions via bootstrapped representations with our devised self-conditioned modeling and self-conditioned guidance, respectively.
@@
  • We perform a systematic study to evaluate the performance of our framework in a variety of real-world and synthetic datasets. The results demonstrate the effectiveness of our framework in comparison to other state-of-the-art baselines.
2 Related Works
2.1 Denoising Diffusion Models for Generation
@@
Recently, denoising diffusion models [45, 15, 46] have been widely adopted in various fields of generation tasks. Their notable capabilities span various domains, including image generation [7, 38, 16], text generation [2, 28, 3], and temporal data modeling [48, 31]. In general, denoising diffusion models are probabilistic models designed to learn a data distribution by denoising variables with normal distributions [47, 25]. In particular, they first create noisy data by progressively adding and intensifying the noise in the clean data, in a Markovian manner. Subsequently, these models learn a denoising network to backtrack each step of the perturbation. During training, the denoising networks are required to predict the clean data or the added noise, given the noisy data after perturbation. After optimization, the denoising networks could be used to generate new data via iterative denoising of noise sampled from a prior distribution [40, 49].
2.2 Graph Generation
@@
Based on the strategies used, graph generation methods could be classified into two categories: (1) One-shot Generation. In this category, the models generate all edges among a defined node set in one single step. One-shot generation models are typically built upon the Variational Autoencoder (VAE) or the Generative Adversarial Network (GAN) structure, aiming to generate edges independently based on the learned latent embeddings. Normalizing flow models [59, 32] propose to estimate the graph density, by establishing an invertible and deterministic function to map latent embeddings to the graphs. More recently, diffusion models have also been adopted for graph generation [26, 21]. To deal with the discrete nature of graph data, DiGress [52] leverages discrete diffusion by considering node and edge types as states in the Markovian transition matrix. (2) Sequential Generation. This strategy entails generating graphs through a series of sequential steps, typically by incrementally adding nodes and connecting them with edges. Models in this category often utilize recurrent networks [29, 57] or Reinforcement Learning (RL) [56] to guide the generation process [43, 1]. Sequential generation is particularly suitable for generating graphs with specific desired properties [60].
2.3 Conditional Generation
@@
Recent works have also explored various strategies to condition generation on specific classes or features. For example, ARM [7] proposes to utilize gradients from classifiers to guide the generation process within each step. LDM [38] enables the incorporation of external information, such as text [37] and semantic maps [20], with a specific encoder to learn the representations. The representations are then incorporated into the underlying UNet backbone [39]. In RCG [27], the authors employ a self-conditioned strategy to condition generation on representations learned from a pre-trained encoder. Despite the advancements in image generation, it still presents significant difficulty when applying these methods in graph generation, due to the complex dataset patterns [4, 26, 17] and discrete sequential generation [53, 34, 43]. In contrast, our framework deals with these challenges with self-conditioned modeling and guidance to enhance graph generation efficacy.
@@
Figure 2: The overall process of GraphRCG during training. Specifically, we learn from the training data a representation generator that outputs a representation based on noise sampled from a standard Gaussian distribution. After that, a graph generator is trained to generate new graphs from noisy graphs under the guidance of the representations.
3 GraphRCG: Self-Conditioned Graph Generation
@@
Our self-conditioned graph generation framework consists of two modules: self-conditioned modeling and self-conditioned guidance for RQ1 and RQ2, respectively. As illustrated in Fig. 2, in self-conditioned modeling, we employ a representation generator to capture graph distributions by learning to denoise representations with noise. The alignment loss between noisy representations and noisy graphs acts as a self-supervised loss to train the encoder. With the optimized representation generator, we train the graph generator by denoising graphs with added noise. At each generation step, we perform self-conditioned guidance via bootstrapped representations with noise at the same timestep. We further enhance guidance by aligning the denoised graph with the clean representations.
@@
3.1 Self-Conditioned Modeling
@@
Representation Generator. Given the representations of all training data provided by the graph encoder, the representation generator is required to learn to generate representations with the same distribution. To enhance generation performance, following [27], we utilize the Representation Diffusion Model (RDM) architecture, which generates representations from Gaussian noise, based on the process of Denoising Diffusion Implicit Models (DDIM) [46]. In particular, RDM utilizes a backbone comprising a fully connected network with multiple residual blocks. Each block is composed of an input layer, a timestep embedding projection layer, and an output layer. The number of residual blocks and the hidden dimension size both act as hyper-parameters. During training, given a representation of a graph sample, learned by the graph encoder, we first perturb it by adding random noise as follows:
$\mathbf{h}_{t}=\sqrt{\alpha_{t}}\mathbf{h}_{0}+\sqrt{1-\alpha_{t}}\epsilon,\quad\text{where}\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
@@
$\mathcal{L}_{RG}=\mathbb{E}_{\mathbf{h}_{0},\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\mathbf{h}_{0}-f_{\gamma}\left(\mathbf{h}_{t},t\right)\right\|_{2}^{2}\right],$ (2)
where $\mathbf{h}_{0}$ is sampled from representations of graphs in training data, and $t$ is uniformly sampled from $\{1,2,\dotsc,T\}$. The target of the representation generator $f_{\gamma}$ is to capture the representation distribution and learn to generate representations from random noise. After optimization, the representation generator could perform sampling from random noise, following the DDIM strategy [46], to obtain new representations.
@@
$\mathcal{L}_{AR}=\mathbb{E}_{G,t}\left[\left\|\mathbf{h}_{t}-h_{\eta}(\mathbf{X}_{t},\mathbf{E}_{t})\right\|_{2}^{2}\right],$ (3)
@@
$q(G_{t}|G_{t-1})=(\mathbf{X}_{t-1}\mathbf{Q}_{t}^{\mathbf{X}},\mathbf{E}_{t-1}\mathbf{Q}_{t}^{\mathbf{E}}),$ (4)
@@
$q(G_{t}|G_{0})=(\mathbf{X}_{0}\prod_{i=1}^{t}\mathbf{Q}_{i}^{\mathbf{X}},\mathbf{E}_{0}\prod_{i=1}^{t}\mathbf{Q}_{i}^{\mathbf{E}}),$ (5)
@@
To specify the Markov transition matrices, previous works have explored several feasible choices. The most prevalent choices in the literature have been uniform transitions [2, 55] and absorbing transitions [4, 24]. However, these do not contain the distribution information and thus could not benefit the capturing of graph distribution in our framework. Therefore, in our approach, we leverage the marginal transitions [18, 52], in which the probability of transitioning to any given state is directly related to its marginal probability observed in the dataset. In this manner, the transition matrices are modeled in a way that mirrors the natural distribution of states in graph data. As the edges are generally sparse in the graph data, the probability of jumping to the state of “no edge” is significantly higher than that of other states. To present the noise process, we first define $\mathbf{p}^{\mathbf{X}}\in\mathbb{R}^{a}$ and $\mathbf{p}^{\mathbf{E}}\in\mathbb{R}^{b}$ as the marginal distributions for the node and edge types, respectively. The marginal transition matrices for nodes and edges are formulated as
$\mathbf{Q}_{t}^{\mathbf{X}}=\alpha^{t}\mathbf{I}+\beta^{t}\mathbf{1}_{a}(\mathbf{p}^{\mathbf{X}})^{\top},\ \ \text{and}\ \ \mathbf{Q}_{t}^{\mathbf{E}}$ $=\alpha^{t}\mathbf{I}+\beta^{t}\mathbf{1}_{b}(\mathbf{p}^{\mathbf{E}})^{\top}.$ (6)
@@
3.2 Self-Conditioned Guidance (Training)
In self-conditioned guidance for RQ2, we optimize a graph generator to create new graphs conditioned on bootstrapped representations from the representation generator. Existing works have explored various methods to perform generation conditioned on specific properties, such as high activity in molecular graph generation [17]. However, these properties are generally several scalar values, which could capture only a small fraction of the information in the dataset. Moreover, these approaches often adopt a classifier or regressor to guide the generation process, which could not exploit the useful information in representations. In contrast, using representations as guidance for generation could largely benefit from the learned distribution in our representation generator.